# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ Abstract.
Data-driven techniques for analysis, modeling, and control of complex dynamical systems are on the uptake. Koopman theory provides the theoretical foundation for the extremely popular kernel extended dynamic mode decomposition (kEDMD). In this work we propose a novel kEDMD scheme to approximate nonlinear control systems accompanied by an in-depth error analysis. The main features of the method are flexible sampling, regularization-based robustness, and an adroit decomposition into micro and macro grids. In addition, we prove proportionality, i.e., explicit dependence on the distance to the (controlled) equilibrium, of the derived uniform bounds on the full approximation error. Leveraging this key property, we rigorously show that asymptotic stability of the data-driven surrogate (control) system implies asymptotic stability of the original (control) system and vice versa.
@@
F.M.P was funded by the Carl Zeiss Foundation within the project DeepTurb–Deep Learning in and from Turbulence. He was further supported by the free state of Thuringia and the German Federal Ministry of Education and Research (BMBF) within the project THInKI–Thüringer Hochschulinitiative für KI im Studium. K.W. gratefully acknowledges funding by the German Research Foundation (DFG) – Project-ID 507037103.
@@ 1. Introduction
Data-driven methods for analysis, modelling, and control of dynamical systems has recently attracted considerable attention, see, e.g., the survey articles [36, 24] and the references therein. A key aspect is the Lyapunov-based stability analysis of dynamical systems as well as for controller design, see, e.g., [2]. For nonlinear systems, however, guarantees for data-driven approaches require a sophisticated analysis [41], which may, e.g., also be conducted in the behavioral setting relying on input-output data [23]. In this work, we provide a framework that allows to infer stability properties of the original system from its data-driven surrogates (and vice versa). Hereby, we leverage the Koopman operator, which provides a theoretically-sound foundation for analysis and control of dynamical systems through the lens of observables, see, e.g., [39, 25] and [38, 31]. In the Koopman approach, the dynamics is lifted to an infinite-dimensional function space of observables. Therein, data-driven surrogate models are approximated using regression. The most prominent approximation technique is extended dynamic mode decomposition (EDMD; [49]), an algorithm that builds upon a finite dictionary of observables and which has been used in a wide range of challenging applications, see, e.g., [52, 55, 56] for some recent examples.
@@
Various works combining the Koopman approach with Lyapunov arguments exist in the literature. For a robust $\mathcal{H}_{\infty}$-approach, we refer to [26, Chapter 2], while Lyapunov-based stabilization of a chemical reactor was the subject of [28]. The authors of [57] present a neural network-based Koopman approach to feedback design via control Lyapunov functions, while Lyapunov functions were approximated in a Koopman-based manner in [5]. However, to rigorously ensure end-to-end guarantees, a theoretically-sound error analysis including, e.g., convergence rates, of data-driven surrogate models serving as approximants is indispensable. Concerning EDMD, convergence to the Koopman operator in the infinite-data limit was proven in [17] (using also infinitely-many observables). Finite-data error bounds for dynamical systems were, to the best of the authors’ knowledge, first provided in [54, 27]. Then, the probabilistic error bounds were further extended to stochastic (control) systems using both i.i.d. and ergodic sampling in [30] and to kernel EDMD in [34], see also [19] for sharp rates in view of spectral approximations and [18] for long-term ergodic predictions via transfer operators. Recently, using kernel EDMD (kEDMD) embedded in suitably-chosen reproducing kernel Hilbert spaces (RKHSs), error bounds in the supremum norm were rigorously shown [20]. Similar estimates have been shown in [53] via an EDMD-variant using Bernstein polynomials. A key finding in the derivation of these uniform error bounds is the invariance of the RKHS under the Koopman operator, see also [8] and [35] for preliminary results. Assuming such uniform error bounds, stabilizing controllers with end-to-end guarantees can be designed, see, e.g., [9, 22, 43], or [4, 51] within model predictive control. Herein, error bounds of EDMD-based surrogates for control systems that proportional to the distance to the set point played a major role. While pointwise and proportional error bounds are fundamental in these works for controller design, there is, to the best of the authors’ knowledge, no rigorous proof of such bound for an existing EDMD variant for control systems. This work closes this gap. In this context, we point out [13], where limitations of linear surrogate models [16, 37] were thoroughly discussed. Hence, for general nonlinear control-affine systems, the bilinear approach proposed in [44, 48], is superior, see, e.g., [33, 9].
@@
The contribution of this work is two-fold. First, we propose a kEDMD-based surrogate model for nonlinear autonomous dynamics and prove that it inherits stability properties of the ground-truth system (and vice-versa). In particular, we show that asymptotic stability of one system implies practical asymptotic stability of the other. Further, we prove the first proportional uniform error bounds on kernel EDMD approximants of the Koopman operator, which certify that the full approximation error decays proportionally w.r.t. the distance to the equilibrium without imposing restrictive assumptions like invariance of some finitely-generated subspace [9]. If a certain compatibility condition linking the decay rates of the stability-certifying Lyapunov function and the proportional error bound holds, we even prove that asymptotic stability is fully preserved. As a second major contribution, we present a novel data-based Koopman approximants using kEDMD, which allows for flexible sampling of state-control pairs and outputs a control-affine surrogate system. We analyze our algorithm and prove proportional bounds on the uniform approximation error between the original model and its surrogate. The proportionality is then leveraged to prove that feedback laws stabilizing the data-driven model also stabilize the original control system (and vice versa). All results are extended in view of regularized kEDMD approximants to improve robustness, e.g., for noisy data.
@@
The outline of the paper is as follows: In Section 2, we begin with a brief recap on Koopman approximants for autonomous systems generated by kernel EDMD. Then, we present uniform bounds on the full approximation error using RKHS interpolation in Section 2.2, before we conclude the section by proposing a data-driven surrogate and the respective error bound. In the subsequent Section 3, we show that (practical) asymptotic stability is inherited from the original dynamical system by the data-driven surrogate and vice versa. In Section 4, we propose a novel and highly-flexible approximation scheme for control systems and rigorously show bounds on the full approximation error. We numerically illustrate and validate our findings in  Section 3.3 and Section 4.3, before conclusions are drawn in Section 5.
[…]
@@ 2. Koopman operator and kernel EDMD
Throughout this paper, let $\Omega\subset\mathbb{R}^{n}$ be an open domain with Lipschitz boundary in the sense of [1, §4.9] [^1] This Lipschitz condition implies the usual cone conditions for interpolation estimates with reproducing kernel Hilbert spaces, cf. [20, Appendix A].. We consider the discrete-time dynamical system given by
@@
with a map $F:\Omega\to\mathbb{R}^{n}$. Here, $x\in\Omega$ and $x^{+}$ are the current and the successor state of the dynamical system Equation DS, respectively. Further, we abbreviate the image of the set $\Omega$ w.r.t. the dynamics $F$ by $\Xi$, i.e., $\Xi=F(\Omega):=\{F(x):x\in\Omega\}$, and assume that $F\in\mathcal{C}^{1}(\Omega,\Xi)$ is a diffeomorphism satisfying the regularity condition $\inf_{x\in\Omega}|\det DF(x)|>0$.
@@
The dynamical system Equation DS can be inferred from the ordinary differential equation
@@
where $x(t)\in\Omega$ represents the state at time $t\geq 0$ and $g_{0}:\Omega\to\mathbb{R}^{n}$ is a locally Lipschitz-continuous map. Then, for a given time step $\Delta t>0$, we can associate the discrete-time dynamical system
$x^{+}=F(\bar{x}):=\bar{x}+\int_{0}^{\Delta t}g_{0}(x(s;\bar{x}))\,\mathrm{d}s,$
using the integral representation of the solution $x(\cdot,\bar{x})$ emanating from the (current) state $\bar{x}\in\Omega$ tacitly assuming that the solution $x(s;\bar{x})$ exists on the time interval $[0,\Delta t]$. We point out that the imposed regularity assumption $\inf_{x\in\Omega}|\det DF(x)|>0$ automatically holds for a sufficiently small time step $\Delta t$.
The linear Koopman operator $\mathcal{K}:C_{b}(\Xi)\to C_{b}(\Omega)$ associated with the system Equation DS maps functions (so-called observables) in $C_{b}(\Xi)$ to functions in $C_{b}(\Omega)$ along the flow of the system and is defined by
@@
or, for short, $\mathcal{K}f=f\circ F$.
@@ 2.1. Kernel EDMD (kEDMD)
the symmetric kernel matrix $(\mathsf{k}(x_{i},x_{j}))_{i,j=1}^{d}$ is positive definite [^2] In the literature, a positive definite kernel is defined by positive semi-definiteness of the kernel matrices. Since we require these matrices to be invertible, we added the term strictly.. The canonical feature $\phi_{z}$ of $\mathsf{k}$ at $z\in\mathbb{R}^{n}$ is defined by $\phi_{z}(x)=\mathsf{k}(z,x)$, $x\in\mathbb{R}^{n}$. It is well known that the linear space $\operatorname{span}\{\phi_{z}:z\in\mathbb{R}^{n}\}$ extends by completion to a Hilbert space $\mathbb{H}$ of functions, the reproducing kernel Hilbert space (RKHS) or simply native space associated with (or generated by) the kernel $\mathsf{k}$, see, e.g., [32, Theorem 2.14]. Continuity of the kernel is inherited by the functions in $\mathbb{H}$ so that $f|_{\Omega}\in C_{b}(\Omega)$ holds for all $f\in\mathbb{H}$. For all functions $f\in\mathbb{H}$, we have the reproducing property
$f(x)=\langle f,\phi_{x}\rangle\qquad\forall\,x\in\mathbb{R}^{n}.$ (2)
We provide an example of popular radially symmetric kernels with compact support.
@@
where $p_{n,k}$ is an univariate polynomial of degree $\left\lfloor\frac{n}{2}\right\rfloor+3k+1$ and $\phi_{n,k}\in\mathcal{C}^{2k}([0,\infty),\mathbb{R})$, see [46, Theorem 9.13]. The Wendland RBF $\Phi_{n,k}$ induces the kernel $\mathsf{k}_{n,k}$ given by $\mathsf{k}_{n,k}(x,y)=\Phi_{n,k}(\|x-y\|)$ for $x,y\in\mathbb{R}^{n}$. By $\mathcal{N}_{\Phi_{n,k}}(\mathcal{O})$, we denote the native space corresponding to the RKHS $\mathbb{H}$ generated by the Wendland kernel on any bounded open domain $\mathcal{O}\subset\mathbb{R}^{n}$ with Lipschitz boundary. Here, denoting by $H^{s}(\mathcal{O})$ the $L^{2}$-Sobolev space of regularity order $s$ on $\mathcal{O}$, we have the identity
@@
with equivalent norms, see e.g. [46, Corollary 10.48] for integer Sobolev orders or [20, Theorem 4.1] for fractional Sobolev orders orders.
To obtain error estimates with finitely many data points, in the following we assume that the set $\Omega$ is bounded. Moreover, we assume forward invariance of $\Omega$ w.r.t. the dynamics Equation DS in order to keep the presentation technically simple. In view of the stability analysis of the subsequent sections, we point out that this seemingly restrictive assumption holds on a suitably chosen sublevel set of the considered Lyapunov function. For a detailed treatment of the case $\Xi=F(\Omega)\not\subset\Omega$, we refer to [20].
A key ingredient to show bounds on the approximation error for kernel-based approximations of the Koopman operator is that the considered RKHS $\mathbb{H}$ is invariant w.r.t. the Koopman operator $\mathcal{K}:C_{b}(\Omega)\to C_{b}(\Omega)$, i.e.,
$\mathcal{K}\,\mathbb{H}\subset\mathbb{H}.$ (4)
This assumption was verified for Wendland kernels (as introduced in Example 2.2) in [20, Section 4.2]. To this end, the authors leverage that these native spaces coincide with fractional Sobolev spaces with equivalent norms, cf. identity Equation 3. We note that the same also applies for Matérn kernels, as illustrated in Remark 2.6. However, we emphasize that assuming the invariance Equation 4 may, in general, lead to severe restrictions, e.g., implying affine-linear dynamics Equation DS as rigorously shown for the RKHS $\mathbb{H}$ generated by Gaussian kernels in [8].
Before introducing kernel-based approximations of the Koopman operator, we briefly note an important consequence of the reproducing property Equation 2 for the Koopman operator: A point $x^{*}\in\Omega$ is an equilibrium of the dynamics Equation DS if and only if $(\mathcal{K}f)(x^{*})=f(x^{*})$ for all $f\in\mathbb{H}$. Indeed, sufficiency is trivial by definition of the Koopman operator. For necessity, assume that $(\mathcal{K}f)(x^{*})=f(x^{*})$ for all $f\in\mathbb{H}$. Then, for all $f\in\mathbb{H}$ we have
$\langle f,\phi_{F(x^{*})}\rangle=f(F(x^{*}))=(\mathcal{K}f)(x^{*})=f(x^{*})=\langle f,\phi_{x^{*}}\rangle$
and, thus, $\phi_{F(x^{*})}=\phi_{x^{*}}$. The strict positive definiteness of the kernel $\mathsf{k}$ then implies $F(x^{*})=x^{*}$, i.e., that $x^{*}$ is an equilibrium of Equation DS.
Next, we briefly recap kernel extended dynamic mode decomposition (kEDMD) as an advanced tool to approximate the Koopman operator from data and refer to [50, 15] and [20, Section 3.2] for further details. For the set of pairwise distinct data points $\mathcal{X}$ given by Equation 1, we set
$V_{\mathcal{X}}=\operatorname{span}\{\phi_{x_{1}},\ldots,\phi_{x_{d}}\},$
where $\phi_{x_{i}}=\mathsf{k}(x_{i},\cdot)$, $i\in[1:d]$. Further, let $P_{\mathcal{X}}$ denote the orthogonal projection in $\mathbb{H}$ onto $V_{\mathcal{X}}$, i.e., for given $f\in\mathbb{H}$, the function $P_{\mathcal{X}}f$ solves the regression problem
@@
cf. [20]. Hence, $g^{*}\in V_{\mathcal{X}}$ is the solution of Equation 5 if and only if $g^{*}=P_{\mathcal{X}}f$. Further, it may be easily seen that Equation 5 is equivalent to $\min_{g\in V_{\mathcal{X}}}\sum_{i=1}^{d}|f(x_{i})-g(x_{i})|^{2}$. Note that $P_{\mathcal{X}}f$ for $f\in\mathbb{H}$ is in fact the interpolation of $f$ at the points in $\mathcal{X}$ in the sense that $P_{\mathcal{X}}f$ is the unique function in $V_{\mathcal{X}}$ which coincides with $f$ at all data points in $\mathcal{X}$. Then, as proven in [20, Proposition 3.2], a kEDMD approximant of the Koopman operator $\mathcal{K}$ on $\mathbb{H}$ is given by
@@
where
$\mathbf{k}_{\mathcal{X}}=[\phi_{x_{1}},\ldots,\phi_{x_{d}}]^{\top}\qquad\text{and}\qquad\mathbf{K}_{\mathcal{X}}=(k(x_{i},x_{j}))_{i,j=1}^{d}.$
In fact, the kEDMD regression solution is $P_{\mathcal{X}}\mathcal{K}|_{V_{\mathcal{X}}}$, which is a linear map from the finite-dimensional space $V_{\mathcal{X}}$ into itself. Correspondingly, kEDMD may also be understood as a method for approximating the Koopman operator by finite-rank operators with range in $V_{\mathcal{X}}$ despite the Koopman operator acting on an infinite-dimensional function space.
We close this subsection by relating the above approximation Equation 6 to another kernel-based surrogate of the Koopman operator.
@@
In [20], another approximant of $\mathcal{K}$ has been defined by $\widetilde{\mathcal{K}}=P_{\mathcal{X}}\mathcal{K}P_{\mathcal{X}}$. The key difference is the following: For the computation of $\widehat{\mathcal{K}}f$, the observable $f\in\mathbb{H}$ has to be propagated by the flow, i.e., we require data samples $(f(x_{i}),f(F(x_{i}))$, $i\in[1:d]$. In contrast, for the computation of the alternative surrogate $\widetilde{\mathcal{K}}f$, the canonical features at the data sites, i.e., $\phi_{x_{i}}$, $i\in[1:d]$, have to be propagated and the observable $f$ only has to be interpolated, i.e., measurements $f(F(x_{i}))$ are not necessary. We refer the interested reader to [20] for a detailed discussion of the pros and cons. Moreover, we provide a collection of results on the surrogate arising from the approximant $\widetilde{\mathcal{K}}$ in Appendix A, which correspond to those that we will derive for $\widehat{\mathcal{K}}$ in the sequel.

### v3

@@ Abstract.
Data-driven techniques for analysis, modeling, and control of complex dynamical systems are on the uptake. Koopman theory provides the theoretical foundation for the popular kernel extended dynamic mode decomposition (kEDMD). In this work, we propose a novel kEDMD scheme to approximate nonlinear control systems accompanied by an in-depth error analysis. Key features are regularization-based robustness and an adroit decomposition into micro and macro grids enabling flexible sampling. But foremost, we prove proportionality, i.e., explicit dependence on the distance to the (controlled) equilibrium, of the derived bound on the full approximation error. Leveraging this key property, we rigorously show that asymptotic stability of the data-driven surrogate (control) system implies asymptotic stability of the original (control) system and vice versa.
@@
F.M.P was funded by the Carl Zeiss Foundation within the project DeepTurb–Deep Learning in and from Turbulence. He was further supported by the free state of Thuringia and the German Federal Ministry of Education and Research (BMBF) within the project THInKI–Thüringer Hochschulinitiative für KI im Studium. L.B. gratefully acknowledges funding by the German Research Foundation (DFG) – Project-ID 545246093.
@@ 1. Introduction
Data-driven methods for analysis, modelling, and control of dynamical systems have recently attracted considerable attention, see, e.g., the survey articles [44, 31] and the references therein. Key aspects are the Lyapunov-based stability analysis of and the controller design for dynamical systems, see, e.g., [2]. For nonlinear systems, however, guarantees for data-driven approaches require more sophisticated techniques [50]. In this work, we provide a framework that allows to infer stability properties of the original system from its data-driven surrogates (and vice versa). Hereby, we leverage the Koopman operator, which provides a theoretically-sound foundation for analysis and control of dynamical systems through the lens of observable functions, see, e.g., [46, 32, 39]. The underlying idea is to lift the dynamics to an infinite-dimensional function space. Then, finite-dimensional surrogate models are constructed using regression in a data-driven fashion. The most prominent approximation scheme is extended dynamic mode decomposition (EDMD; [59]), an algorithm that builds upon a finite dictionary of observables and has been used in a wide range of challenging applications, see, e.g., [61, 64] for some recent examples.
@@
Various works combining the Koopman approach with Lyapunov arguments exist in the literature. For a robust $\mathcal{H}_{\infty}$-approach, we refer to [33, Chapter 2], while Lyapunov-based stabilization of a chemical reactor was the subject of [35]. In [65], a neural-network-based Koopman approach to feedback design via control Lyapunov functions is presented, while Lyapunov functions were approximated in a Koopman-based manner in [6]. However, to rigorously ensure end-to-end guarantees, an error analysis including, e.g., convergence rates, of data-driven surrogate models serving as approximants is indispensable. Concerning EDMD, strong convergence to the Koopman operator in the infinite-data limit was proven in [24] (using also infinitely-many observables). Finite-data error bounds for dynamical systems were, to the best of the authors’ knowledge, first provided in [34, 38, 63]. Then, the probabilistic error bounds [34, 63] were further extended to stochastic (control) systems using both i.i.d. and ergodic sampling in [38] and to kernel EDMD in [42], see also [27] for sharp rates in view of spectral approximations and [26] for long-term ergodic predictions via transfer operators. Recently, using kernel EDMD (kEDMD) embedded in suitably-chosen reproducing kernel Hilbert spaces (RKHSs), error bounds in the supremum norm were rigorously shown in [28]. Similar estimates have been shown in [62] via an EDMD-variant using Bernstein polynomials. A key finding in the derivation of these uniform error bounds is the invariance of the RKHS under the Koopman operator, see also [13] and [43] for preliminary results. Assuming such uniform error bounds, stabilizing controllers with end-to-end guarantees can be designed, see, e.g., [14, 30, 52], or [4, 60] within model predictive control. Herein, error bounds on EDMD-based surrogates for control systems that are proportional to the distance to the set point play a major role. While pointwise and proportional error bounds are fundamental in these works for controller design, there is, to the best of the authors’ knowledge, no rigorous proof of such bound for an existing EDMD variant for control systems. This work closes this gap. In this context, we point out [19], where limitations of linear surrogate models [25, 45] were thoroughly discussed. Hence, for general nonlinear control-affine systems, the bilinear approach proposed in [53, 58], is superior, see, e.g., [41, 14].
@@
The contribution of this work is two-fold. First, we propose a kEDMD-based surrogate model for nonlinear autonomous dynamics and prove that it inherits stability properties of the ground-truth system (and vice-versa). In particular, we show that asymptotic stability of one system implies practical asymptotic stability of the other. Further, we prove the first uniform proportional error bounds on kernel EDMD approximants of the Koopman operator, which certify that the full approximation error decays proportionally w.r.t. the distance to the equilibrium without imposing restrictive assumptions like invariance of some finitely-generated subspace [14]. If a certain compatibility condition linking the decay rates of the stability-certifying Lyapunov function and the proportional error bound holds, we even prove that asymptotic stability is fully preserved. As a second major contribution, we present novel data-based Koopman approximants using kEDMD, which allow for flexible sampling of state-control pairs and yield control-affine surrogate systems. We analyze our algorithm and prove proportional bounds on the uniform approximation error between the original model and its surrogate. The proportionality is then leveraged to prove that feedback laws stabilizing the data-driven model also stabilize the original control system (and vice versa). All results are extended in view of regularized kEDMD approximants to improve robustness, e.g., for noisy data.
@@
The outline of the paper is as follows: In Section 2, we begin with a brief recap on Koopman approximants for autonomous systems generated by kernel EDMD. Then, we present uniform bounds on the full approximation error using RKHS interpolation, before we conclude the section by proposing a data-driven surrogate and the respective error bound. In the subsequent Section 3, we show that (practical) asymptotic stability is inherited from the original dynamical system by the data-driven surrogate and vice versa. In Section 4, we propose a novel and highly-flexible approximation scheme for control systems, prove error bounds on the full approximation error and illustrate the results numerically. Conclusions are drawn in Section 5.
Notation. We use $\|\cdot\|$ for the Euclidean norm on $\mathbb{R}^{n}$ and its induced matrix norm on $\mathbb{R}^{n\times n}$. The Frobenius norm on $\mathbb{R}^{n\times m}$ will be denoted by $\|\cdot\|_{F}$. Moreover, we let $\mathbb{R}_{\geq 0}=[0,\infty)$. Further, we set $\mathbb{N}=\mathbb{Z}_{\geq 1}$, and for $d\in\mathbb{N}$ we use the abbreviation $1:d]:=\mathbb{Z}\cap[1,d]$. By $C_{b}(\Omega)$, we denote the space of bounded continuous functions on a set $\Omega\subset\mathbb{R}^{n}$. Moduli of continuity of a continuous function $f:\Omega\to\mathbb{R}$ are denoted by $\omega_{f}$. Recall that a modulus of continuity $\omega_{f}:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ vanishes at zero, is continuous at zero, and satisfies $|f(x)-f(y)|\leq\omega_{f}(\|x-y\|)\qquad\forall\,x,y\in\Omega$. We define comparison functions to introduce our stability notions analogously to [15], see, e.g., [22]. A continuous, strictly increasing function $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ satisfying $\alpha(0)=0$ is said to be of class $\mathscr{K}$. If it is, in addition, unbounded, the comparison function $\alpha$ is class $\mathscr{K}_{\infty}$. A continuous function $\beta:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}_{\geq 0}$ is called a class $\mathscr{K\!L}$-function if, for each $t\geq 0$, $\beta(\cdot,t)\in\mathscr{K}_{\infty}$ and $\beta(r,\cdot)$ is strictly monotonically decreasing with $\lim_{t\rightarrow\infty}\beta(r,t)=0$ for all $r>0$.
@@ 2. Koopman operator and kernel EDMD
Throughout this paper, let $\Omega\subset\mathbb{R}^{n}$ be an open domain with Lipschitz boundary in the sense of [1, §4.9]. [^1] The imposed Lipschitz condition implies the usual cone conditions for interpolation estimates with reproducing kernel Hilbert spaces, cf. [28, Appendix A]. We consider the discrete-time dynamical system given by
@@
with a map $F:\Omega\to\mathbb{R}^{n}$. Here, $x\in\Omega$ and $x^{+}$ are the current and the successor state of the dynamical system eq. DS, respectively. Further, we abbreviate the image of the set $\Omega$ w.r.t. the dynamics $F$ by $\Xi$, i.e., $\Xi=F(\Omega):=\{F(x):x\in\Omega\}$, and assume that $F\in\mathcal{C}^{1}(\Omega,\Xi)$ is a diffeomorphism satisfying the regularity condition $\inf_{x\in\Omega}|\det DF(x)|>0$.
@@
The dynamical system eq. DS can be inferred from the ordinary differential equation
@@
where $x(t)\in\Omega$ represents the state at time $t\geq 0$ and $g_{0}:\Omega\to\mathbb{R}^{n}$ is a locally Lipschitz-continuous map. Then, for a given time step $\Delta t>0$, we can associate the discrete-time dynamical system $x^{+}=F(\bar{x}):=\bar{x}+\int_{0}^{\Delta t}g_{0}(x(s;\bar{x}))\,\mathrm{d}s$, using the integral representation of the solution $x(\cdot,\bar{x})$ emanating from the (current) state $\bar{x}\in\Omega$ tacitly assuming that the solution $x(s;\bar{x})$ exists on the time interval $0,\Delta t]$. We point out that the imposed regularity assumption $\inf_{x\in\Omega}|\det DF(x)|>0$ automatically holds for a sufficiently small time step $\Delta t$.
The linear Koopman operator $\mathcal{K}:C_{b}(\Xi)\to C_{b}(\Omega)$ associated with the system eq. DS maps functions (so-called observables) in $C_{b}(\Xi)$ to functions in $C_{b}(\Omega)$ along the flow of the system and is defined by
@@
or, for short, $\mathcal{K}f=f\circ F$. In the following, we assume forward invariance of $\Omega$ w.r.t. the dynamics eq. DS in order to keep the presentation technically simple. In view of the stability analysis of the subsequent sections, we point out that this seemingly restrictive assumption holds on a suitably chosen sublevel set of the considered Lyapunov function. For a detailed treatment of the case $\Xi=F(\Omega)\not\subset\Omega$, we refer to [28].
@@ 2.1. Kernel EDMD (kEDMD)
the symmetric kernel matrix $(\mathsf{k}(x_{i},x_{j}))_{i,j=1}^{d}$ is positive definite. [^2] In the literature, a positive definite kernel is defined by positive semi-definiteness of the kernel matrices. Since we require these matrices to be invertible, we added the term strictly. The canonical feature $\phi_{z}$ of $\mathsf{k}$ at $z\in\mathbb{R}^{n}$ is defined by $\phi_{z}(x)=\mathsf{k}(z,x)$, $x\in\mathbb{R}^{n}$. It is well known that the linear space $\operatorname{span}\{\phi_{z}:z\in\mathbb{R}^{n}\}$ extends by completion to a Hilbert space $\mathbb{H}$ of functions, the reproducing kernel Hilbert space (RKHS) or simply native space associated with (or generated by) the kernel $\mathsf{k}$, see, e.g., [40, Theorem 2.14]. Continuity of the kernel is inherited by the functions in $\mathbb{H}$ so that $f|_{\Omega}\in C_{b}(\Omega)$ holds for all $f\in\mathbb{H}$. For all functions $f\in\mathbb{H}$, we have the reproducing property
$f(x)=\langle f,\phi_{x}\rangle\qquad\forall\,x\in\mathbb{R}^{n},$ (2)
where $\langle\cdot,\cdot\rangle$ is the inner product of the RKHS $\mathbb{H}$. Before we proceed, we provide an example of popular radially symmetric kernels with compact support.
@@
where $p_{n,k}$ is a univariate polynomial of degree $\left\lfloor\frac{n}{2}\right\rfloor+3k+1$ and $\phi_{n,k}\in\mathcal{C}^{2k}([0,\infty),\mathbb{R})$, see [55, Theorem 9.13]. The Wendland RBF $\Phi_{n,k}$ induces the kernel $\mathsf{k}_{n,k}$ given by $\mathsf{k}_{n,k}(x,y)=\Phi_{n,k}(\|x-y\|)$ for $x,y\in\mathbb{R}^{n}$. By $\mathcal{N}_{\Phi_{n,k}}(\mathcal{O})$, we denote the native space corresponding to the RKHS $\mathbb{H}$ generated by the Wendland kernel on any bounded open domain $\mathcal{O}\subset\mathbb{R}^{n}$ with Lipschitz boundary. Here, denoting by $H^{s}(\mathcal{O})$ the $L^{2}$-Sobolev space of regularity order $s$ on $\mathcal{O}$, we have the identity
@@
with equivalent norms, see e.g. [55, Corollary 10.48] for integer Sobolev orders or [28, Theorem 4.1] for fractional Sobolev orders orders.
To obtain error estimates with finitely many data points, in the following we assume that the set $\Omega$ is bounded. Moreover, we choose the kernel $\mathsf{k}$ such that the corresponding RKHS $\mathbb{H}$ is invariant w.r.t. the action of the Koopman operator $\mathcal{K}:C_{b}(\Omega)\to C_{b}(\Omega)$, i.e., $\mathcal{K}\,\mathbb{H}\subset\mathbb{H}$, a key property to derive bounds on the approximation error for kernel-based approximations of the Koopman operator. A suitable choice are Wendland kernels (see example 2.2) as rigorously shown in [28, Section 4.2]. The authors of [28] leverage that these native spaces coincide with fractional Sobolev spaces with equivalent norms, cf. identity eq. 3. We note that the same holds true for Matérn kernels, as illustrated in remark 2.6. We point out that, in general, the invariance $\mathcal{K}\,\mathbb{H}\subset\mathbb{H}$ does not hold for nonlinear systems, e.g., the RKHS $\mathbb{H}$ generated by Gaussian kernels only satisfies this invariance condition if the dynamics (DS) are affine linear as shown in [13].
Before introducing kernel-based approximations of the Koopman operator, we briefly note an important consequence of the reproducing property eq. 2 for the Koopman operator: A point $x^{*}\in\Omega$ is an equilibrium of the dynamics eq. DS if and only if $(\mathcal{K}f)(x^{*})=f(x^{*})$ for all $f\in\mathbb{H}$. Indeed, necessity is trivial by definition of the Koopman operator. For sufficiency, assume that $(\mathcal{K}f)(x^{*})=f(x^{*})$ for all $f\in\mathbb{H}$. Then, for all $f\in\mathbb{H}$ we have $\langle f,\phi_{F(x^{*})}\rangle=f(F(x^{*}))=(\mathcal{K}f)(x^{*})=f(x^{*})=\langle f,\phi_{x^{*}}\rangle$ and, thus, $\phi_{F(x^{*})}=\phi_{x^{*}}$. The strict positive definiteness of the kernel $\mathsf{k}$ then implies $F(x^{*})=x^{*}$, i.e., that $x^{*}$ is an equilibrium of eq. DS.
Next, we briefly recap kernel extended dynamic mode decomposition (kEDMD) as an advanced tool to approximate the Koopman operator from data and refer to [57, 23] and [28, Section 3.2] for further details. For the set of pairwise distinct data points $\mathcal{X}$ given by eq. 1, we set $V_{\mathcal{X}}=\operatorname{span}\{\phi_{x_{1}},\ldots,\phi_{x_{d}}\}$, where $\phi_{x_{i}}=\mathsf{k}(x_{i},\cdot)$, $i\in[1:d]$. Further, let $P_{\mathcal{X}}$ denote the orthogonal projection in $\mathbb{H}$ onto $V_{\mathcal{X}}$, i.e., for given $f\in\mathbb{H}$, the function $P_{\mathcal{X}}f$ solves the regression problem
@@
cf. [28]. Hence, $g^{*}\in V_{\mathcal{X}}$ is the solution of eq. 4 if and only if $g^{*}=P_{\mathcal{X}}f$. Further, it may be easily seen that eq. 4 is equivalent to $\min_{g\in V_{\mathcal{X}}}\sum_{i=1}^{d}|f(x_{i})-g(x_{i})|^{2}$. Note that $P_{\mathcal{X}}f$ for $f\in\mathbb{H}$ is in fact the interpolation of $f$ at the points in $\mathcal{X}$ in the sense that $P_{\mathcal{X}}f$ is the unique function in $V_{\mathcal{X}}$ which coincides with $f$ at all data points in $\mathcal{X}$. Then, as proven in [28, Proposition 3.2], a kEDMD approximant of the Koopman operator $\mathcal{K}$ on $\mathbb{H}$ is given by
@@
where $\mathbf{k}_{\mathcal{X}}=[\phi_{x_{1}},\ldots,\phi_{x_{d}}]^{\top}$ and $\mathbf{K}_{\mathcal{X}}=(k(x_{i},x_{j}))_{i,j=1}^{d}$. In fact, the kEDMD regression solution is $P_{\mathcal{X}}\mathcal{K}|_{V_{\mathcal{X}}}$, which is a linear map from the finite-dimensional space $V_{\mathcal{X}}$ into itself. Correspondingly, kEDMD may also be understood as a method for approximating the Koopman operator by finite-rank operators with range in $V_{\mathcal{X}}$ despite the Koopman operator acting on an infinite-dimensional function space.
We close this subsection by relating the above approximation eq. 5 to another kernel-based surrogate of the Koopman operator.
@@
In [28], another approximant of $\mathcal{K}$ has been defined by $\widetilde{\mathcal{K}}=P_{\mathcal{X}}\mathcal{K}P_{\mathcal{X}}$. The key difference is the following: For the computation of $\widehat{\mathcal{K}}f$, the observable $f\in\mathbb{H}$ has to be propagated by the flow, i.e., we require data samples $(f(x_{i}),f(F(x_{i}))$, $i\in[1:d]$. In contrast, for the computation of the alternative surrogate $\widetilde{\mathcal{K}}f$, the canonical features $\phi_{x_{i}}$ at the data sites have to be propagated, while the observable $f$ only has to be interpolated, i.e., measurements $f(F(x_{i}))$ are not necessary. We refer the interested reader to [28] for a detailed discussion. Moreover, a collection of results on the surrogate arising from the approximant $\widetilde{\mathcal{K}}$ – which correspond to those that we will derive for $\widehat{\mathcal{K}}$ – can be found in the (additional) appendix, see Appendix A.